What is the probability of selecting a red card and then an ace without replacement from a standard deck of cards?
Answer by stanbon(75887)

What is the probability of selecting a red card and then an ace without replacement from a standard deck of cards?
-----
P(ace | red card) = P(red AND ace)/P(red) = (2/52)/(1/2) = 4/52 = 1/13
================
Cheers,
stan H.
============